OpenStudy (anonymous):

I need to know how to set up the following problem. I can work it out, but am having trouble with the initial set up. A tank and pail contain a total of 5,136 ml of water. Jacob pours 314 ml of water from the pail into the tank. The amount of water in the tank is now seven times what is left in the pail. How much water was in the pail at first?? Please help. Need to help my son with this problem to turn in tomorrow.

OpenStudy (anonymous):

Let the initial amount of water in the tank be x. Let the initial amount of water in the pail be y. x + y = 5136 x +314 = 7(y - 314) Now that you have two equations and two unknown variables, you can solve for x and y.

OpenStudy (anonymous):

x + 314 = 7y - 2198 x + 314 = AAAGGHH!! Still confused! What am I doing wrong???

OpenStudy (anonymous):

seems I am having more difficulty with this one that i anticipated :-(

OpenStudy (anonymous):

That's alright! When you have two equations and two unknowns, you solve for x and y like this: Take one equation and solve for one of the variables: I'll use: x + y =5186 and solve for x. x = 5186 - y

OpenStudy (anonymous):

Then take that x, and plug it in for the x in the second equation. x + 314 = 7(y - 314) becomes... (5186 - y) + 314 = 7(y - 314)

OpenStudy (anonymous):

Now move the terms around and solve for y. 5186 - y + 314 = 7y - 2198 5500 - y = 7y - 2198 8y = 7698 y = 962.25 mL ---> this is the initial amount of water in the pail

OpenStudy (anonymous):

Check me to make sure my algebra is right. I did it a little fast.

OpenStudy (anonymous):

i was working it too.. u were faster than me, but we were the same up to 8y... how did u get 8?

OpenStudy (anonymous):

you have to collect all the "y" terms on the same side. If move the "-y" on the left to the right side with the "7y" you get 8y

OpenStudy (anonymous):

i really really appreciate u help

OpenStudy (anonymous):

still trying figure it out :-)

OpenStudy (anonymous):

no problem. let me know if you still are having trouble with it

OpenStudy (anonymous):

thank you again... really appreciate it